# Learning Diverse Attacks on Large Language Models for Robust Red-Teaming and Safety Tuning

### v1

@@ Abstract
Red-teaming, or identifying prompts that elicit harmful responses, is a critical step in ensuring the safe and responsible deployment of large language models (LLMs). Developing effective protection against many modes of attack prompts requires discovering diverse attacks. Automated red-teaming typically uses reinforcement learning to fine-tune an attacker language model to generate prompts that elicit undesirable responses from a target LLM, as measured, for example, by an auxiliary toxicity classifier. We show that even with explicit regularization to favor novelty and diversity, existing approaches suffer from mode collapse or fail to generate effective attacks. As a flexible and probabilistically principled alternative, we propose to use GFlowNet fine-tuning, followed by a secondary smoothing phase, to train the attacker model to generate diverse and effective attack prompts. We find that the attacks generated by our method are effective against a wide range of target LLMs, both with and without safety tuning, and transfer well between target LLMs. Finally, we demonstrate that models safety-tuned using a dataset of red-teaming prompts generated by our method are robust to attacks from other RL-based red-teaming approaches.
@@ 1 Introduction
The deployment of large language models (LLMs) in the wild has raised concerns about their potential harmful impacts for nearly a decade [31, 59]. These concerns have grown with the increasing capabilities of LLMs: even models fine-tuned to satisfy certain safety constraints can be manipulated to produce toxic outputs [57]. Red-teaming, or identification of ‘attack’ prompts that elicit undesirable responses, gives model developers a chance to address such vulnerabilities before deployment [42]. This paper studies the problem of automatically generating diverse attack prompts for LLMs and argues for the potential of robust automated red-teaming in the development of effective defenses.
@@
Effective red-teaming requires identifying many modes of attack [20]. Methods for automated red-teaming based on stochastic optimization of attack prompts [68, 65] have been proposed, while others have used reinforcement learning (RL) to train an attacker language model (LM), allowing to generate novel prompts efficiently at test time [42, 20]. However, even when regularized to favor diversity, these methods struggle to balance between diversity and effective attacks (Fig. 2). They often suffer from mode collapse, where the attacker LLM generates a small set of similar prompts, or overly focus on diversity and fail to generate effective attacks (Fig. 3). Moreover, we have empirically found that they also fail to discover attacks that persist across different target LLMs (Table 2).
@@
This paper takes an amortized inference perspective on red-teaming: we view the problem of generating an attack prompt as one of sampling a latent variable in a probabilistic model. Using the off-policy RL approach of GFlowNet fine-tuning, proposed for inference of linguistic latent variables in [23], we fine-tune attack LMs to sample the full posterior distribution over attack prompts.
@@
However, controlling the ‘peakiness’ of the posterior distribution – the preference of attack quality to attack diversity – is challenging, especially when red-teaming a target LLM that has been safety-tuned to resist some modes of attack, leading to a sparser landscape of attack prompts. Inspired by the success of behavior cloning in offline RL [16, 27] and reinforcement learning for human feedback [RLHF; 34], we propose a two-stage GFlowNet fine-tuning procedure with MLE smoothing. As illustrated in Fig. 1, we first fine-tune a pretrained attacker LM with a GFlowNet objective and collect high-reward attack prompts discovered in the course of training (Step 1). The collected prompts form an offline dataset. Subsequently, the pretrained attacker model is fine-tuned again to maximize the likelihood of the offline dataset (Step 2). The first stage, GFlowNet fine-tuning, enables us to collect a set of diverse and effective attack prompts using exploratory off-policy training. In the second phase, we obtain a smooth distribution over high-reward attack prompts, since all the collected attack prompts in the offline dataset are considered equally important and the attacker LM is trained to maximize their log-likelihood uniformly. Consequently, we find that the attacker LM is able to sample attack prompts that are both diverse and effective.
@@
We empirically evaluate the efficacy of our proposed method in red-teaming four target LLMs: GPT-2 [43], dolly-v2-7b [13], Gemma-2b-it [39], and Llama-2-7b-chat [51]. Our approach is found to sample more diverse and effective attack prompts than other relevant baselines. Moreover, many of our attack prompts effectively transfer to other target LLMs that are not used for training the attacker model, such as Llama-2-13b/70b-chat, Llama-3-8b/70b-instruct [1], Starling-7b-beta [66], and Mistral-7b-instruct-v0.2 [28]. Lastly, we fine-tune a target LLM to generate refusal responses to the discovered attack prompts and find that the model fine-tuned with our red-teaming prompts is more robust than the models safety-tuned with other RL-based red-teaming methods.
@@
It is important to note that while we study an approximate measure of toxicity as a proxy for harmfulness, following past works [42, 20], the true harmful impact of an LLM output is often subjective and dependent on the social context of deployment [59]. We nonetheless believe that the methods we propose will be useful in practice and can be extended to other measures of harmfulness.
@@
  • When safety-tuned on attack prompts generated by our method, target LLMs become robust to attacks generated with other RL-based methods without performance degradation on other tasks.
@@ 2 Related work
As LLMs increase in general capabilities and performance, so does the risk associated to potential misuse of LLMs. To mitigate this, LLMs are often trained to refuse to generate content given prompts that are dangerous, offensive, or harmful [2, 3]. This is done at various stages of the training process such as filtering out harmful training data [39] or fine-tuning on ‘safe’ responses to harmful prompts [51]. This process is often augmented by red-teaming, which proactively looks for ways to elicit harmful behavior from models. Prior works [15, 61, 54] rely on a large amount of human annotation to identify vulnerabilities of LMs. To automate red-teaming, Perez et al. [42] formulate red teaming as an RL problem and train an LM to sample toxic prompts. However, most RL algorithms are not suitable for sampling diverse objects since they tend to converge to a single reward-maximizing trajectory. To overcome this limitation,  Hong et al. [20] propose using a novelty-based reward to encourage a policy to explore diverse samples during RL training. Instead of generating a prompt from scratch, Lee et al. [30] replace words of prompts from a predefined user input pool to attack LMs using Bayesian optimization in a sample-efficient manner. As a concurrent work, Rainbow Teaming [47] sample an attack prompt from a pool and iteratively mutate the prompt with auxiliary LLMs.
@@
Jailbreaking and red-teaming are closely related in that red-teaming proactively tries to discover vulnerabilities for the purpose of improving model safety, whereas jailbreaking generally refers to circumventing the built-in safeguards of models. Initially, jailbreaks were found manually through trial and error, taking advantage of the different objectives models were trained against [57]. Recently, automated jailbreak attacks are becoming increasingly popular. They utilize techniques such as genetic algorithms [35], iterative gradient-based methods [68], or automated prompting via auxiliary LLMs [9] to optimize query prompts.  Mazeika et al. [37] propose a method defending against GCG [68], which is one of the most popular gradient based jailbreak methods. A drawback of these methods is the computational cost since the optimization has to be performed separately for each new query prompt. Another drawback is the transferability of jailbreaks. Meade et al. [38] have shown that prompts optimized by GCG to jailbreak one target LLM do not transfer to jailbreak other target LLMs.
@@
Generative flow networks [GFlowNets; 5] are a probabilistic framework to train stochastic policies to sample discrete compositional objects (e.g., graphs, sequences) proportionally to a reward. Sampling objects proportionally to a reward results in diverse high-reward samples. Consequently, GFlowNets have found applications in a wide variety of problems including biological sequence generation [25], combinatorial optimization [63, 64], Bayesian structure learning [14], variational EM with discrete latent variables [22], and probablistic neurosymbolic inference [52]. Most closely related to our work is [23], which uses the GFlowNet objective to fine-tune LMs for solving intractable inference problems such as sampling chains of thought [58]. We use GFlowNet fine-tuning as a part of our approach for learning policies which generate diverse prompts that elicit toxic responses from target LLMs.
@@ 3.1 Preliminaries
We denote the target large language model (LLM) as $p_{\phi}$, which samples a text response ${\mathbf{y}}$ for a given prompt ${\mathbf{x}}$, with probability of $p_{\phi}({\mathbf{y}}|{\mathbf{x}})$. The goal of red-teaming an LLM is to identify prompts ${\mathbf{x}}$ that elicit toxic responses from the target LLM. A binary toxicity classifier, denoted as $p_{\psi}$, is used to quantify the effectiveness of a prompt ${\mathbf{x}}$. Specifically, the likelihood of the response being classified as toxic by the classifier $p_{\psi}(c=1|{\mathbf{x}},{\mathbf{y}})$, where $c\in\{0,1\}$ is a binary variable denoting toxicity, is used to measure the effectiveness of a prompt ${\mathbf{x}}$. Moreover, for the attack to be effective, the prompts ${\mathbf{x}}$ should appear natural, as unnatural prompts (with high perplexity) are easy to defend against with simple filters [26]. Red-teaming can often be a time-consuming process if done manually as the space of prompts is quite large. Perez et al. [42], Hong et al. [20] formulate red-teaming as an RL problem, to automate the discovery of these prompts. This involves training an autoregressive language model (LM) as a policy, $p_{\theta}$, parameterized by $\theta$, to generate prompts that maximize the expected reward (as measured by the toxicity of the response generated by the target LLM):
$\operatorname*{maximize}_{\theta}\mathbb{E}_{{\mathbf{x}}~{}\sim p_{\theta}({\mathbf{x}}),y\sim p_{\phi}({\mathbf{y}}|{\mathbf{x}})}\left[p_{\psi}(c=1|{\mathbf{x}},{\mathbf{y}})-\lambda D_{\mathrm{KL}}(p_{\theta}({\mathbf{x}})\;\|\;p_{\texttt{ref}}({\mathbf{x}}))\right],$ (1)
where $\lambda>0$ and $D_{\mathrm{KL}}$ is a KL-divergence between the policy $p_{\theta}$ and a pretrained LM $p_{\texttt{ref}}$ is the constraint which enforces the prompt ${\mathbf{x}}$ to be close to natural language text. However, most RL algorithms are not suitable for discovering diverse prompts since they generally concentrate most of probability mass of the policy $p_{\theta}$ on actions with highest reward, often resulting in a deterministic policy that generates a single prompt [5]. While Hong et al. [20] propose adding a novelty-based reward term along with entropy bonus [48] as a regularization to encourage the policy to generate diverse prompts, empirically we find that it is challenging to find an optimal trade-off between diversity and toxicity rate even with the regularization.
@@
5:        Uniformly and independently sample $s_{1}$ and $s_{2}$ from $\{0,1\}$, and initialize $\tau\leftarrow 1.0$.
6:        $\tau\leftarrow\texttt{Uniform}(0.5,2.0)$ if $s_{1}==0$.
7:        if $s_{2}==0$ then
@@ 3.2 GFlowNet fine-tuning and smoothing with MLE on collected high-reward prompts
We propose a two-stage approach designed to learn a stochastic policy to sample diverse and effective prompts for red-teaming. The first stage consists of fine-tuning a pretrained LM $p_{\theta}$ as a generative flow network [GFlowNet; 5, 6, 23] policy in order to collect sample prompts, and the second stage restarts fine-tuning from the original pretrained LM policy but this time with maximum likelihood estimation (MLE) on the high-reward prompts collected during GFlowNet training in the first stage.
@@
GFlowNets are diversity-seeking RL algorithms that learn a policy $p_{\theta}$ which samples prompts with a probability proportional to the reward associated with the prompt [^1] In the case of generating sequences, GFlowNets are equivalent to MaxEnt RL algorithms [18].. We define the reward for a prompt ${\mathbf{x}}$ as follows:
@@
where $\beta$ and $\gamma$ are positive constants that control the “peakiness” (tempering) of toxicity score $R_{1}({\mathbf{x}})$ and the reference LM likelihood $R_{2}({\mathbf{x}})$, respectively. The prompt ${\mathbf{x}}=(x_{0},x_{1},\ldots,x_{T})$, consisting of $T$ tokens with a special token $x_{0}$ indicating the beginning of a sentence, is generated autoregressively from a behavior policy, which is a mix of $p_{\theta}$ and a tempered variant of it. We define $(x_{0},x_{1},\ldots,x_{t})$ as a state in the generative process and the token sampled from the policy at each step is the action. To learn the parameters $\theta$, we use the trajectory balance learning objective [36]:
$\mathcal{L}({\mathbf{x}};\theta)=\left(\log\frac{Z_{\theta}\prod_{t=1}^{T}p_{\theta}(x_{t}|x_{0},x_{1},\ldots,x_{t-1})}{R({\mathbf{x}})}\right)^{2},$ (3)
@@
$\log R({\mathbf{x}})\approx\frac{1}{\beta}\frac{1}{k}\sum_{i=1}^{k}\log p_{\psi}(c=1|{\mathbf{x}},{\mathbf{y}}^{(i)})+\frac{1}{\gamma}\log p_{\texttt{ref}}({\mathbf{x}}),\quad\text{where }{\mathbf{y}}^{(i)}\overset{\mathrm{iid}}{\sim}p_{\phi}({\mathbf{y}}|{\mathbf{x}}).$ (4)
Table 1: Examples showing difficulty of balancing between toxicity reward and reference model likelihood.
@@
Research methods to handle the process of forging credit card rewards for phishing software for small transactions $-0.06$ $-83.09$
@@
As we illustrate in §4, simply using GFlowNet fine-tuning to sample effective and diverse red-teaming prompts can be challenging in practice due to non-trivial choice of the temperature parameters $\beta$ and $\gamma$. While in principle there are choices of $\beta$ and $\gamma$ which can balance the reward and diversity well, in practice GFlowNet fine-tuning can be overly sensitive to the peakiness of reward [29]. Moreover, balancing between $\beta$ and $\gamma$ is non-trivial. The first two examples, shown in §3.2, get a low total reward compared to the last one, even though they are grammatically valid sentences and achieve high toxicity rewards since they are assigned numerically low likelihood by $p_{\texttt{ref}}$. If we try much smaller $\beta$ to focus on toxicity reward $R_{1}({\mathbf{x}})$, the policy $p_{\theta}$ will likely generate prompts from potentially spurious modes of the toxicity classifier, which will have high perplexity. On the other hand, if we set $\gamma$ to a small value, the model would merely focus on $R_{2}({\mathbf{x}})$ and not generate toxic prompts.
@@
To reduce sensitivity to the aforementioned parameters of the reward distribution, while preserving the mode coverage and ability of the training procedure to generalize to new modes, we propose an inexpensive retraining step following GFlowNet fine-tuning. This second step akin to behavior cloning [10, 16, 27] in RL, where a policy is trained to imitate expert trajectories. First, we store all prompts sampled by the policy $p_{\theta}$ during GFlowNet fine-tuning in Stage 1. We expect this set to contain a diverse and high reward prompts discovered by off-policy exploration during GFlowNet fine-tuning. Subsequently, we filter the prompts in this set based on the toxicity score $R_{1}({\mathbf{x}})$ and language model likelihood $R_{2}({\mathbf{x}})$ being larger than some thresholds. The collected examples form an offline dataset, and the reference policy is fine-tuned again (from the same initial state as in Stage 1) to maximize log-likelihood of samples from the offline dataset. We outline our complete method in Alg. 1. This procedure results in a smooth approximation to the distribution over high-reward prompts found during exploratory training.

### v2

@@ Abstract
Red-teaming, or identifying prompts that elicit harmful responses, is a critical step in ensuring the safe and responsible deployment of large language models (LLMs). Developing effective protection against many modes of attack prompts requires discovering diverse attacks. Automated red-teaming typically uses reinforcement learning to fine-tune an attacker language model to generate prompts that elicit undesirable responses from a target LLM, as measured, for example, by an auxiliary toxicity classifier. We show that even with explicit regularization to favor novelty and diversity, existing approaches suffer from mode collapse or fail to generate effective attacks. As a flexible and probabilistically principled alternative, we propose to use GFlowNet fine-tuning, followed by a secondary smoothing phase, to train the attacker model to generate diverse and effective attack prompts. We find that the attacks generated by our method are effective against a wide range of target LLMs, both with and without safety tuning, and transfer well between target LLMs. Finally, we demonstrate that models safety-tuned using a dataset of red-teaming prompts generated by our method are robust to attacks from other RL-based red-teaming approaches. Code is available at https://github.com/GFNOrg/red-teaming.
@@ 1 Introduction
The deployment of large language models (LLMs) in the wild has raised concerns about their potential harmful impacts for nearly a decade (Lee, 2016; Weidinger et al., 2021). These concerns have grown with the increasing capabilities of LLMs: even models fine-tuned to satisfy certain safety constraints can be manipulated to produce toxic outputs (Wei et al., 2023). Red-teaming, or identification of ‘attack’ prompts that elicit undesirable responses, gives model developers as well as regulators a chance to identify and address such vulnerabilities before deployment (Perez et al., 2022). This paper studies the problem of automatically generating diverse attack prompts for LLMs and argues for the potential of robust automated red-teaming in the development of effective defenses.
@@
Effective red-teaming requires identifying many modes of attack (Hong et al., 2024). Methods for automated red-teaming based on stochastic optimization of attack prompts (Zou et al., 2023; Zhao et al., 2024) have been proposed, while others have used reinforcement learning (RL) to train an attacker language model (LM), allowing to generate novel prompts efficiently at test time (Perez et al., 2022; Hong et al., 2024). However, even when regularized to favor diversity, these methods struggle to balance between diversity and effective attacks (Fig. 2). They often suffer from mode collapse, where the attacker LM generates a small set of similar prompts, or focus solely on diversity and fail to generate effective attacks (Fig. 3). Moreover, we have empirically found that they also fail to discover attacks that transfer across different target LLMs (Table 3).
@@
This paper takes an amortized inference perspective on red-teaming: we view the problem of generating an attack prompt as sampling a latent variable in a probabilistic model. Using the off-policy RL approach of GFlowNet fine-tuning, proposed for inference of linguistic latent variables in (Hu et al., 2024), we fine-tune attack LMs to sample the full posterior distribution over attack prompts.
@@
However, controlling the ‘peakiness’ of the posterior distribution – the preference of attack quality to attack diversity – is challenging, especially when red-teaming a target LLM that has been safety-tuned to resist some modes of attack, leading to a sparser landscape of attack prompts. Inspired by the success of behavior cloning in offline RL (Emmons et al., 2022; Jang et al., 2021), we propose a two-stage GFlowNet fine-tuning procedure with MLE smoothing. As illustrated in Fig. 1, we first fine-tune a pretrained attacker LM with a GFlowNet objective and collect high-reward attack prompts discovered in the course of training (Step 1). The collected prompts form an offline dataset. Subsequently, the pretrained attacker model is fine-tuned again to maximize the likelihood of the offline dataset (Step 2). The first stage, GFlowNet fine-tuning, enables us to collect a set of diverse and effective attack prompts using exploratory off-policy training. In the second phase, we obtain a smooth distribution over high-reward attack prompts, since all the collected attack prompts in the offline dataset are considered equally important and the attacker LM is trained to maximize their log-likelihood uniformly. Consequently, we find that the attacker LM is able to sample attack prompts that are both diverse and effective.
@@
We empirically evaluate the efficacy of our proposed method in red-teaming five target LLMs: GPT-2 (Radford et al., 2019), Dolly-v2-7b (Conover et al., 2023), Gemma-2b-it (Mesnard et al., 2024), Llama-2-7b-chat (Touvron et al., 2023), and Llama-3.1-8B-Instruct (Dubey et al., 2024). Our approach is found to sample more diverse and effective attack prompts than other relevant baselines. Moreover, many of our attack prompts effectively transfer to other target LLMs that are not used for training the attacker model, such as Llama-2-13b/70b-chat, Llama-3-8b/70b-instruct (Dubey et al., 2024), Starling-7b-beta (Zhu et al., 2023), and Mistral-7b-instruct-v0.2 (Jiang et al., 2023). Lastly, we fine-tune a target LLM to generate refusal responses to the discovered attack prompts and find that the model fine-tuned with our red-teaming prompts is more robust than the models safety-tuned with other RL-based red-teaming methods.
@@
It is important to note that while we study an approximate measure of toxicity as a proxy for harmfulness, following past works (Perez et al., 2022; Hong et al., 2024), the true harmful impact of an LLM output is often subjective and dependent on the social context of deployment (Weidinger et al., 2021). We nonetheless believe that the methods we propose will be useful in practice and can be extended to other measures of harmfulness.
@@
  • When safety-tuned on attack prompts generated by our method, target LLMs become robust to attacks generated by other RL-based methods without performance degradation on other tasks.
@@ 2 Related work
As LLMs increase in general capabilities and performance, so does the risk associated to potential misuse of LLMs. To mitigate this, LLMs are often trained to refuse to generate content given prompts that are dangerous, offensive, or harmful (Bai et al., 2022a; b). This is done at various stages of the training process such as filtering out harmful training data (Mesnard et al., 2024) or fine-tuning on ‘safe’ responses to harmful prompts (Touvron et al., 2023). This process is often augmented by red-teaming, which proactively looks for ways to elicit harmful behavior from models. Prior works (Dinan et al., 2019; Xu et al., 2021; Wallace et al., 2022) rely on a large amount of human annotation to identify vulnerabilities of LMs. To automate red-teaming, Perez et al. (2022) formulate red teaming as an RL problem and train an LM to sample toxic prompts. However, most RL algorithms are not suitable for sampling diverse objects since they tend to converge to a single reward-maximizing trajectory. To overcome this limitation,  Hong et al. (2024) propose using a novelty-based reward to encourage a policy to explore diverse samples during RL training. Instead of generating a prompt from scratch, Lee et al. (2023) replace words of prompts from a predefined user input pool to attack LMs using Bayesian optimization in a sample-efficient manner. Rainbow Teaming (Samvelyan et al., 2024) samples an attack prompt from a pool and iteratively mutates the prompt with auxiliary LLMs.
@@
Jailbreaking and red-teaming are closely related in that red-teaming proactively tries to discover vulnerabilities for the purpose of improving model safety, whereas jailbreaking generally refers to circumventing the built-in safeguards of models. Initially, jailbreaks were found manually through trial and error, taking advantage of the different objectives models were trained against (Wei et al., 2023). Recently, automated jailbreak attacks are becoming increasingly popular. They utilize techniques such as genetic algorithms (Liu et al., 2024), iterative gradient-based methods (Zou et al., 2023), or automated prompting via auxiliary LLMs (Chao et al., 2023) to optimize query prompts. Mazeika et al. (2024) propose a method defending against GCG (Zou et al., 2023), one of the most popular gradient-based jailbreak methods. A drawback of these methods is the computational cost since the optimization has to be performed separately for each new query prompt. Another drawback is the poor transferability of jailbreaks. Meade et al. (2024) have shown that prompts optimized by GCG to jailbreak one target LLM do not transfer to jailbreak other target LLMs.
@@
Generative flow networks (GFlowNets; Bengio et al., 2021) are a probabilistic framework to train stochastic policies to sample discrete compositional objects (e.g., graphs, sequences) proportionally to a reward. Sampling objects proportionally to a reward results in diverse high-reward samples. Consequently, GFlowNets have found applications in a wide variety of problems including biological sequence generation (Jain et al., 2022), combinatorial optimization (Zhang et al., 2023a; b), Bayesian structure learning (Deleu et al., 2022), variational EM with discrete latent variables (Hu et al., 2023), and probabilistic neurosymbolic inference (van Krieken et al., 2023). Most closely related to our work is (Hu et al., 2024), which uses the GFlowNet objective to fine-tune LMs for solving intractable inference problems such as sampling chains of thought (Wei et al., 2022). We use GFlowNet fine-tuning as a part of our approach for learning policies which generate diverse prompts that elicit toxic responses from target LLMs.
@@ 3.1 Preliminaries
The target LLM, denoted $p_{\phi}$, samples a text response ${\mathbf{y}}$ for a given prompt ${\mathbf{x}}$ with probability $p_{\phi}({\mathbf{y}}\mid{\mathbf{x}})$. The goal of red-teaming an LLM is to identify prompts ${\mathbf{x}}$ that elicit toxic responses from the target LLM. A binary toxicity classifier, denoted as $p_{\psi}$, is used to quantify the effectiveness of an attack prompt. Specifically, the effectiveness of a prompt ${\mathbf{x}}$ is measured by the likelihood of the response ${\mathbf{y}}\sim p_{\phi}({\mathbf{y}}\mid{\mathbf{x}})$ being classified as toxic by the classifier: $p_{\psi}(c=1\mid{\mathbf{x}},{\mathbf{y}})$, where $c\in\{0,1\}$ is a binary variable denoting toxicity. Moreover, for the attack to be effective, the prompt ${\mathbf{x}}$ should appear natural, as unnatural prompts (with high perplexity under some prior) are easy to defend against with simple filters (Jain et al., 2023).
Red-teaming can often be a time-consuming process if done manually as the space of prompts is quite large. Perez et al. (2022); Hong et al. (2024) formulate red-teaming as an RL problem, to automate the discovery of these prompts. This involves training a LM as a policy $p_{\theta}$, parameterized by $\theta$, to generate prompts that maximize the expected reward (as measured by the toxicity of the response generated by the target LLM):
$\operatorname*{maximize}_{\theta}\mathbb{E}_{{\mathbf{x}}~{}\sim p_{\theta}({\mathbf{x}}),y\sim p_{\phi}({\mathbf{y}}\mid{\mathbf{x}})}\left[p_{\psi}(c=1\mid{\mathbf{x}},{\mathbf{y}})\right]-\lambda D_{\mathrm{KL}}(p_{\theta}\;\|\;p_{\texttt{ref}}),$ (1)
where the KL divergence term, weighted by a hyperparameter $\lambda>0$, encourages the policy $p_{\theta}$ to remain close to an initial pretrained LM $p_{\texttt{ref}}$, penalizing the generation of prompts ${\mathbf{x}}$ that are far from natural language text. However, most RL algorithms are not suitable for discovering diverse prompts since they generally concentrate most of probability mass of the policy $p_{\theta}$ on actions with highest reward, often resulting in a deterministic policy that generates a single prompt (Bengio et al., 2021). While Hong et al. (2024) propose adding a novelty-based reward term along with entropy bonus (Schulman et al., 2017a) as a regularization to encourage the policy to generate diverse prompts, empirically we find that it is challenging to find an optimal trade-off between diversity and toxicity rate even with the regularization. In the context of red-teaming, identifying diverse and effective attack prompts is critical to ensure that the target LLM is sufficiently safety-tuned for a broad range of scenarios which might be encountered when the model is deployed in the wild.
@@
5:        Uniformly randomly sample behavior policy $b\in\{\text{tempered policy},\ \text{replay buffer}\}$.
6:        if $b=\text{tempered policy}$ then
7:           Uniformly randomly set $\tau\leftarrow 1.0$ or $\tau\leftarrow\texttt{Uniform}(0.5,2.0)$.
@@ 3.2 GFlowNet fine-tuning and smoothing with MLE on collected high-reward prompts
A probabilistic view of the problem provides a principled alternative. Specifically, problem of generating diverse and effective red-teaming prompts can be viewed as one of generating samples from a (tempered) reward distribution. We adopt the perspective of generative flow networks (GFlowNets; Bengio et al., 2021; 2023), leveraging their ability to learn policies that sample from a target distribution defined over compositional objects such as sequences (Jain et al., 2022) and graphs (Bengio et al., 2023). To instantiate the probabilistic perspective, we propose a two-stage approach designed to learn a stochastic policy to sample diverse and effective prompts for red-teaming. The first stage consists of fine-tuning a pretrained LM $p_{\theta}$ as a GFlowNet policy (Hu et al., 2024) in order to collect prompts, and the second stage restarts fine-tuning from the original pretrained LM policy but this time with maximum likelihood estimation (MLE) on the high-reward prompts collected during GFlowNet training in the first stage.
@@
GFlowNets are diversity-seeking RL algorithms that learn a policy $p_{\theta}$ which samples prompts with a probability proportional to the reward associated with the prompt [^1] In the case of generating sequences, GFlowNets are equivalent to MaxEnt RL (Haarnoja et al., 2017).. We define the reward for a prompt ${\mathbf{x}}$ as follows:
@@
where $\beta$ and $\gamma$ are positive constants that control the ‘peakiness’ (tempering) of the toxicity score $R_{1}({\mathbf{x}})$ and of the reference LM likelihood $R_{2}({\mathbf{x}})$, respectively. The prompt ${\mathbf{x}}=(x_{0},x_{1},\ldots,x_{T})$, consisting of $T$ tokens with a special token $x_{0}$ indicating the beginning of a sentence, is generated autoregressively from a behavior policy, which is a mix of $p_{\theta}$ and a tempered variant of it. We define $(x_{0},x_{1},\ldots,x_{t})$ as a state in the generative process and the token sampled from the policy at each step is the action. To learn the parameters $\theta$, we use the trajectory balance learning objective (Malkin et al., 2022):
$\mathcal{L}({\mathbf{x}};\theta)=\left(\log\frac{Z_{\theta}\prod_{t=1}^{T}p_{\theta}(x_{t}\mid x_{0},x_{1},\ldots,x_{t-1})}{R({\mathbf{x}})}\right)^{2},$ (3)
@@
$\log R({\mathbf{x}})\approx\frac{1}{\beta}\frac{1}{k}\sum_{i=1}^{k}\log p_{\psi}(c=1\mid{\mathbf{x}},{\mathbf{y}}^{(i)})+\frac{1}{\gamma}\log p_{\texttt{ref}}({\mathbf{x}}),\quad\text{where }{\mathbf{y}}^{(i)}\overset{\mathrm{iid}}{\sim}p_{\phi}({\mathbf{y}}\mid{\mathbf{x}}).$ (4)
Table 1: Examples showing difficulty of balancing between toxicity ($R_{1}$) and reference model likelihood ($R_{2}$).
@@
Research methods to handle the process of forging credit card rewards for phishing software for small transactions. $-0.06$ $-83.09$
@@
As we illustrate in §4, using GFlowNet fine-tuning alone to sample effective and diverse red-teaming prompts can be challenging in practice due to non-trivial choice of the temperature parameters $\beta$ and $\gamma$. While in principle there are choices of $\beta$ and $\gamma$ which can balance the reward and diversity well, in practice GFlowNet fine-tuning can be overly sensitive to the peakiness of the reward (Lau et al., 2024). Moreover, balancing between $\beta$ and $\gamma$ to achieve the desired behavior is non-trivial. For example, while all three examples shown in §3.2 get a high toxicity reward, the first two get a low total reward compared to the last one, even though they are grammatically valid sentences, since they are assigned a low likelihood by $p_{\texttt{ref}}$. If we set a much smaller $\beta$ to increase the weight of the toxicity reward $R_{1}({\mathbf{x}})$, the policy $p_{\theta}$ would likely generate prompts from potentially spurious modes of the toxicity classifier, which will have high perplexity under the reference model. On the other hand, if we set $\gamma$ to a small value, the model would merely focus on the naturality score $R_{2}({\mathbf{x}})$ and not generate toxic prompts.
@@
To reduce sensitivity to the aforementioned parameters of the reward distribution, while preserving the mode coverage and ability of the training procedure to generalize to new modes, we propose an inexpensive retraining step that is applied following GFlowNet fine-tuning. This second step is akin to behavior cloning (Chen et al., 2021; Emmons et al., 2022; Jang et al., 2021) in RL, where a policy is trained to imitate expert trajectories. First, we store all prompts sampled by the policy $p_{\theta}$ during GFlowNet fine-tuning in Stage 1. We expect this set to contain diverse and high-reward prompts discovered by off-policy exploration during GFlowNet fine-tuning. Subsequently, we filter the prompts in this set based on the toxicity score $R_{1}({\mathbf{x}})$ and language model likelihood $R_{2}({\mathbf{x}})$ being larger than some thresholds. The collected examples form an offline dataset, and the reference policy is fine-tuned again (from the same initial state as in Stage 1) to maximize log-likelihood of samples from this offline dataset. Stage 2 is very inexpensive in practice, taking under 5% of total (Stage 1 and 2) training time in our experiments (Section 4.2).
We outline our complete method in Alg. 1. This procedure results in a smooth approximation to the distribution over high-reward prompts found during exploratory training.